# Interpretable Clustering: A Survey

### v1

@@ Abstract
In recent years, much of the research on clustering algorithms has primarily focused on enhancing their accuracy and efficiency, frequently at the expense of interpretability. However, as these methods are increasingly being applied in high-stakes domains such as healthcare, finance, and autonomous systems, the need for transparent and interpretable clustering outcomes has become a critical concern. This is not only necessary for gaining user trust but also for satisfying the growing ethical and regulatory demands in these fields. Ensuring that decisions derived from clustering algorithms can be clearly understood and justified is now a fundamental requirement. To address this need, this paper provides a comprehensive and structured review of the current state of explainable clustering algorithms, identifying key criteria to distinguish between various methods. These insights can effectively assist researchers in making informed decisions about the most suitable explainable clustering methods for specific application contexts, while also promoting the development and adoption of clustering algorithms that are both efficient and transparent.
@@ 1 Introduction
For decades, numerous algorithms have been proposed to solve clustering problems across different applications, achieving high accuracy. However, in most cases, clustering models exist as black boxes, leading to common questions such as: How are the clustering results formed? Can people understand the logic behind the formation of the clustering results? Is the model trustworthy? The clustering model’s ability to explain such issues is tentatively defined as model’s clustering interpretability or explainability [3]. Given that most researchers in data mining and machine learning use interpretability and explainability interchangeably, this paper will use the term interpretability throughout this paper .
To date, interpretability still lacks a precise or mathematical definition. Different sources provide slightly varying definitions——for instance, it is defined as “the ability to explain or to present in understandable terms to a human” in [4], “the degree to which a human can understand the cause of a decision” in [5], and “make the behavior and predictions of machine learning systems understandable to humans” in [6]. Collectively, these definitions can all capture the essence of interpretability.
@@
As far as we know, there have been several reviews that summarize methods related to interpretability. However, these reviews either do not focus on the clustering domain [7, 8, 9, 10, 11] or were published too early to include the latest research [12]. To fill this gap, we have comprehensively collected existing interpretable clustering methods and proposed a set of criteria to classify them, ensuring that all methods related to interpretable clustering can be categorized under one of these criteria. Furthermore, we divide the clustering process into three stages and classify all interpretable clustering methods according to their interpretability at different stages, providing the overall framework for this review: (1) the feature selection stage (pre-clustering), (2) the model building stage (in-clustering), and (3) the model explanation stage (post-clustering). We believe this review will provide readers with a new understanding of interpretable clustering and lay a foundation for future research in this area.
The rest of this paper is organized as follows. Section 2 discusses the need for interpretable clustering. Section 3 provides a taxonomy of interpretable clustering methods. Sections 4 to 6 review interpretable pre-clustering, in-clustering, and post-clustering methods, respectively, based on different stages of interpretability in the clustering process. Finally, Section 7 concludes the paper and discusses future directions.
@@ 2.1 What is interpretable clustering?
Any technology or tool that enhances interpretability in clustering analysis can be categorized under the domain of interpretable clustering. A hallmark of these methods is the integration of interpretable models [14] at any stage of the clustering pipeline. These interpretable elements accompany the final clustering results, making them understandable, trustworthy, and usable by humans. Such elements may include, but are not limited to, the use of specific feature values (e.g., age, income) within the data to identify key factors that contribute to the clustering outcomes. End users can rely on this information to comprehend the clustering results and assess whether the conclusions drawn from them are trustworthy.
@@ 2.2 What is a good interpretable clustering method?
First, the form of interpretable evidence should be as simple as possible. For instance, the number of feature values used to derive a cluster should be minimized, which greatly reduces the complexity for end users in understanding the results. Second, each cluster should contain unique and distinguishable information compared to other clusters. In other words, the same interpretable evidence should ideally lead to one specific cluster without overlapping with others. This uniqueness enhances the credibility of the evidence, ensuring that end users can trust it is closely tied to the specific cluster, thereby reducing confusion with other clusters serving different functions.
To determine the goodness of an interpretable clustering method, or even to quantify it, one must consider the specific interpretable model being used. For example, when utilizing decision tree models, it is clear that the evidence used to define each cluster is highly distinctive through the tree’s splits, thereby satisfying the basic requirement of uniqueness. Additionally, one can measure how easily end users understand the results by examining the structural parameters of the tree [15], such as the number of leaf nodes (i.e., the number of clusters) and the average depth of the tree. The process from data to clusters is represented by paths from the root to the leaf nodes, with each branching node recording the decision (splitting feature value) that leads to a cluster. Using fewer feature values results in more concise interpretable evidence, making it easier for end users to understand and trust the clustering results.
@@ 3 A taxonomy of interpretable clustering methods
In this section, after collecting and summarizing existing interpretable clustering methods, we establish the following criteria to taxonomize them systematically:
Firstly, based on widely recognized clustering processes, existing interpretable clustering methods can be categorized into three types: pre-clustering methods, in-clustering methods, and post-clustering methods. Specifically, pre-clustering methods are typically executed before the clustering process and often relate to the selection of interpretable features. In-clustering methods construct interpretable clustering models for the samples, producing accurate partitions without the need for additional operations. Post-clustering methods, on the other hand, typically focus on interpreting the results of existing clustering models, attempting to explain the outcomes generated by black-box models through interpretable models.
Secondly, most methods, particularly in-clustering and post-clustering methods, can be distinguished by the different interpretable models (as shown in Fig. 1) they utilize, including the following categories:
@@
  • Decision tree: the decision tree model is widely recognized as an interpretable model in machine learning and is commonly used for classification and regression tasks. Its interpretability stems from the recursive, hierarchical splitting of data based on feature values to generate intermediate results, with the final output traceable through the feature values used in the splits. Instances are allocated to different leaf nodes (clusters) determined by specific splitting points according to certain criteria, following a clear, transparent path from the root node (representing the whole dataset) down through the branch nodes, which is easily understood by end users.
  • Rules: unlike decision tree-based models, where the end-user needs to understand how a cluster is derived from the entire dataset by following a hierarchical path through the tree, which becomes progressively intricate as the tree deepens, rule-based methods provide a more direct way of understanding how a cluster is extracted. Interpretability in rule-based methods arises from the generation of candidate rules based on feature values, typically expressed as logical combinations of values at the same level (e.g., meaningful patterns), which are more straightforward for end-users to grasp.
  • Prototype: the concept of a prototype (also referred to as an exemplar) can be understood similarly to the concept of a centroid in the $k$-means algorithm. Each prototype serves as a representative of its corresponding cluster, and samples that are sufficiently close to a given prototype are considered members of that cluster. Meanwhile, it is generally permissible for the samples represented by different prototypes to overlap.
  • Convex polyhedral: this type of interpretable model essentially extends convex polygons from two-dimensional space into higher dimensions, where each cluster is enclosed by a set of bounding planes. Each polyhedron is formed by the intersection of a limited number of half-spaces, effectively defining the boundaries of the clusters in the higher-dimensional space.
  • Description: a description can be defined as a concise and interpretable representation of key features or attributes that characterize a specific concept. For example, in the context of community analysis, a description of a community would outline the distinguishing features of that community, such as shared demographics, behaviors, or attributes, effectively summarizing the community’s internal structure and distinguishing it from other communities.
@@
Thirdly, existing methods can be categorized into model-level and feature-level interpretability based on their degree of explainability. While most of the methods discussed in this paper focus on designing interpretable models to obtain clustering results or fitting the results of third-party algorithms, some methods also emphasize the extraction of interpretable features from complex data, or the investigation of the relationships between specific clusters and their associated features, thereby enhancing interpretability.
Finally, methods can additionally be classified based on the nature of the data they are intended to process. These data types may include tabular data (numeric, categorical, or a combination of both), sequential data (such as discrete sequences and time series), as well as image, text, and graph data.
The taxonomy outlined in Fig. 2 provides a framework for classifying clustering methods according to four distinct criteria. These criteria serve as dimensions through which existing interpretable clustering methods can be comprehensively characterized. At the same time, they can also be employed to identify methods that meet specific interpretability and performance requirements.

### v2

@@ Abstract
In recent years, much of the research on clustering algorithms has primarily focused on enhancing their accuracy and efficiency, frequently at the expense of interpretability. However, as these methods are increasingly being applied in high-stakes domains such as healthcare, finance, and autonomous systems, the need for transparent and interpretable clustering outcomes has become a critical concern. This is not only necessary for gaining user trust but also for satisfying the growing ethical and regulatory demands in these fields. Ensuring that decisions derived from clustering algorithms can be clearly understood and justified is now a fundamental requirement. To address this need, this paper provides a comprehensive and structured review of the current state of explainable clustering algorithms, identifying key criteria to distinguish between various methods. These insights can effectively assist researchers in making informed decisions about the most suitable explainable clustering methods for specific application contexts, while also promoting the development and adoption of clustering algorithms that are both efficient and transparent. For convenient access and reference, an open repository organizes representative and emerging interpretable clustering methods under the taxonomy proposed in this survey, available at https://github.com/hulianyu/Awesome-Interpretable-Clustering
@@ 1 Introduction
For decades, numerous algorithms have been proposed to solve clustering problems across different applications, achieving high accuracy. However, in most cases, clustering models exist as black boxes, leading to common questions such as: How are the clustering results formed? Can people understand the logic behind the formation of the clustering results? Is the model trustworthy? The clustering model’s ability to explain such issues is tentatively defined as model’s clustering interpretability or explainability [3]. Explainability is commonly defined as the extent to which the internal mechanics of a machine learning system can be clarified in human terms, i.e., explanations should be faithful to the model while providing information that is relevant in the current context [4]. Given that most researchers in data mining and machine learning use interpretability and explainability interchangeably [5], this paper will use the term interpretability throughout this paper.
To date, interpretability still lacks a precise or mathematical definition [6, 7, 8]. Different sources provide slightly varying definitions – for instance, it is defined as “the ability to explain or to present in understandable terms to a human” in [9], “the degree to which a human can understand the cause of a decision” in [10], and “make the behavior and predictions of machine learning systems understandable to humans” in [4]. Collectively, these definitions can all capture the essence of interpretability.
@@
As far as we know, there have been several reviews that summarize methods related to interpretability. However, most existing reviews are either too general [11, 12, 13, 14] or focus on specific domains [15, 16, 17, 18], and do not focus on the clustering domain. Moreover, the only survey closely related to this topic [19] was published relatively early and thus does not include the latest research or emerging ideas in this field. To fill this gap, we have comprehensively collected existing interpretable clustering methods and proposed a set of criteria to classify them, ensuring that all methods related to interpretable clustering can be categorized under one of these criteria. Furthermore, we divide the clustering process into three stages and classify all interpretable clustering methods according to their interpretability at different stages, providing the overall framework for this review: (1) the feature selection stage (pre-clustering), (2) the model building stage (in-clustering), and (3) the model explanation stage (post-clustering). We believe this review will provide readers with a new understanding of interpretable clustering and lay a foundation for future research in this area.
The rest of this paper is organized as follows. Section 2 discusses the need of interpretable clustering. Section 3 provides a taxonomy of interpretable clustering methods. Sections 5 to 7 review interpretable pre-clustering, in-clustering, and post-clustering methods, respectively, based on different stages of interpretability in the clustering process. Finally, Section 9 concludes the paper and discusses future directions.
@@ 2.1 What is interpretable clustering?
For instance, in a loan evaluation scenario, a conventional clustering algorithm may integrate all applicant attributes such as age, gender, income, and credit score into its objective function, treating them as equally important in forming clusters. While this may produce well-separated applicant groups, it remains unclear which attributes actually drive these distinctions. In contrast, an interpretable clustering method can reveal the decisive factors such as low credit scores or high debt-to-income ratios that define a cluster, allowing the reasoning behind the grouping to be clearly understood and validated.
Interpretable clustering, in general, aims to make such reasoning explicit by incorporating mechanisms that expose how cluster assignments are determined. Any approach that provides interpretability during the clustering process or makes the resulting outcomes easier to understand can be categorized under this domain. A hallmark of these methods is the integration of interpretable models [21] at various stages of the clustering pipeline. These interpretable components accompany the final clustering results, making them understandable, trustworthy, and usable by humans. Such components may include, but are not limited to, the use of specific feature values (e.g., age, income) combined with explicit model syntax (e.g., rules, trees) to identify key factors that contribute to the clustering outcomes, allowing end users to comprehend the results and validate the conclusions derived from these interpretable elements.
@@ 2.2 What is a good interpretable clustering method?
A good interpretable clustering method can therefore be characterized by the following key properties, reflecting its ability to provide explanations in two complementary aspects:
  • Simplicity and parsimony. The form of interpretable evidence should be as concise as possible, minimizing the number of feature values or conditions used to define each cluster. A simpler explanatory form not only reduces the cognitive load on end users but also allows them to follow the reasoning process more easily. In other words, the fewer features required to describe how a cluster is formed, the more transparent and accessible the explanation becomes.
  • Uniqueness and exclusivity. Each cluster should convey distinct and non-overlapping information, reflecting its own functional role within the data. Ideally, the same interpretable evidence should correspond to only one specific cluster, ensuring that the explanation remains both credible and unambiguous. This exclusivity enhances trust, as end users can be confident that the evidence is tightly linked to a single, well-defined cluster, rather than ambiguously shared among multiple clusters. Overlapping explanations across clusters would blur these distinctions and risk confusion or misinterpretation.
These principles can further guide how the quality of interpretability is evaluated in practice. To determine the goodness of an interpretable clustering method, or even to quantify it, one must consider the specific interpretable model being used. For example, when utilizing decision tree models, it is clear that the evidence used to define each cluster is highly distinctive through the tree’s splits, thereby satisfying the basic requirement of uniqueness. Additionally, one can measure how easily end users understand the results by examining the structural parameters of the tree [22], such as the number of leaf nodes (i.e., the number of clusters) and the average depth of the tree. The process from data to clusters is represented by paths from the root to the leaf nodes, with each branching node recording the decision (splitting feature value) that leads to a cluster. Using fewer feature values results in more concise interpretable evidence, making it easier for end users to understand and trust the clustering results.
@@ 3 A taxonomy of interpretable clustering methods
In this section, after collecting and systematically reviewing existing interpretable clustering methods, we establish a unified taxonomy that organizes them along four distinct criteria, as illustrated in Fig. 1. This taxonomy provides a structured framework for classifying interpretable clustering methods according to four complementary dimensions: (1) Process stage, (2) Interpretable model, (3) Interpretability level, and (4) Data modality. Together, these dimensions comprehensively characterize how existing approaches achieve interpretability and serve as a foundation for identifying appropriate methods under different analytical contexts and requirements.
(1) Process stage. Based on widely recognized clustering processes, interpretable clustering methods can be categorized into three types: pre-clustering, in-clustering, and post-clustering. Pre-clustering methods are typically executed before the clustering process and focus on selecting or extracting interpretable features, aiming to provide human-understandable representations for subsequent clustering. In-clustering methods integrate interpretability directly into the model construction process, generating accurate partitions whose formation mechanisms are inherently transparent, without the need for additional post-hoc explanations. Post-clustering methods, in contrast, interpret the results of existing black-box clustering models by constructing surrogate interpretable models that reveal the reasoning behind previously opaque clustering outcomes.
(2) Interpretable model. A large proportion of in-clustering and post-clustering approaches can be distinguished by the type of interpretable model they employ. As illustrated in Fig. 2, four representative model families are commonly used to construct explainable cluster structures, each offering a distinct mechanism for generating human-understandable evidence of how clusters are formed.
@@
  • Decision tree. The decision tree model is one of the most widely recognized interpretable models in machine learning and has long been used for classification and regression tasks. Its interpretability stems from the recursive, hierarchical splitting of data based on feature values to generate intermediate results, where the final outputs can be traced through the sequence of splits. Instances are allocated to different leaf nodes (clusters) according to specific splitting criteria, following a transparent path from the root node, which represents the entire dataset, down through the branching nodes. This hierarchical structure allows end users to understand both why and how each sample is assigned to a cluster, making the reasoning process explicit and verifiable.
  • Rules. In contrast to decision tree-based models, where users must trace a hierarchical path from the root to a leaf node, rule-based methods provide a more direct way to understand how clusters are defined. Interpretability arises from generating a set of candidate rules based on feature values, typically expressed as logical combinations (conjunctions or disjunctions) of conditions at the same level, such as “age >> 40 and income << 30k.” These flat, non-hierarchical structures describe cluster-defining conditions in a way that is both concise and intuitive for end users, avoiding the progressive complexity that can arise in deep trees.
  • Prototype. The concept of a prototype, also referred to as an exemplar, can be understood analogously to the centroid in the $k$-means algorithm. Each prototype serves as a representative of its corresponding cluster, and samples that are sufficiently close to a given prototype are considered members of that cluster. Interpretability in this case arises from identifying which representative instance or exemplar contributes most to defining the cluster. Unlike hierarchical or rule-based models, prototype-based approaches emphasize representativeness rather than logical structure, and may allow a certain degree of overlap among the clusters being explained.
  • Convex polyhedral. A convex polyhedral model generalizes the geometric intuition of convex polygons to higher dimensions, where each cluster is represented as a convex region enclosed by bounding planes formed by the intersection of a finite number of half-spaces. Interpretability arises from the explicit geometric constraints that delineate each region, allowing cluster membership to be visually and spatially verified. Unlike prototype-based models with boundaries implicitly defined by similarity measures, convex polyhedral models provide explicit and well-defined geometric boundaries.
  • Description-based. Beyond the four canonical forms, data-type-dependent interpretable models exist that do not fit neatly into a single category. A representative example is the description-based approach, where clusters are defined by concise, human-readable predicates or summaries derived directly from data semantics. Unlike decision tree or rule-based models with fixed syntactic forms, or geometric models such as prototype- and polyhedral-based methods that rely on spatial continuity, description-based models are particularly common in relational or graph-structured settings. In community analysis, for instance, explanations often rely on node attributes, structural connectivity, or semantic relations, providing non-vector, structurally grounded summaries that capture each community’s distinctive role and clearly distinguish it from other communities.
Illustrative example. Consider a two-dimensional dataset with two features $(F_{1},F_{2})$ and three natural clusters, as illustrated in Fig. 2. This illustrative example is used to show how different interpretable models provide transparent reasoning about the formation of clusters.
  • Decision tree. The decision tree partitions the feature space through a sequence of hierarchical splits. For example, the model may first divide the data based on whether $F_{2}<2.5$, forming Cluster 1 for samples that satisfy this condition (left branch). For the remaining samples with $F_{2}\geq 2.5$, a second split on $F_{1}>3$ further separates them into Cluster 2 and Cluster 3 ($F_{1}\leq 3$). Each cluster can thus be traced through an explicit decision path from root to leaf. For instance, the path $F_{2}\geq 2.5$ (at the root node) $\rightarrow$ $F_{1}>3$ (at the right branch)” corresponds to Cluster 2, clearly illustrating the sequential reasoning behind the assignment.
  • Rule. Rule-based models describe clusters through a set of flat logical conditions rather than hierarchical decision paths. In this example, the three clusters can be expressed as parallel rules operating at the same logical level: samples satisfying $F_{1}<2.5$ and $F_{2}<2.5$ form Cluster 1; those with $F_{1}>3$ and $F_{2}>2$ form Cluster 2; and those with $F_{1}<2.5$ and $F_{2}>3$ form Cluster 3. Each rule provides a self-contained, human-readable statement specifying the feature-value combinations that define cluster membership. Compared with the decision tree, rule-based models establish direct logical relations between features and clusters.
  • Prototype. Prototype-based models define each cluster through a representative exemplar. In this example, three prototypes $\mathbf{p}_{1}$, $\mathbf{p}_{2}$, and $\mathbf{p}_{3}$ represent the three clusters, each with its own influence region in the feature space. For instance, samples near $\mathbf{p}_{1}$ at $(F_{1}=1,F_{2}=1)$ form Cluster 1, those around $\mathbf{p}_{2}$ at $(F_{1}=4.5,F_{2}=4.5)$ form Cluster 2, and those surrounding $\mathbf{p}_{3}$ at $(F_{1}=1,F_{2}=5)$ form Cluster 3. Interpretability arises from how each prototype locally captures its cluster while all prototypes collectively outline the global cluster structure. Unlike hierarchical or rule-based models relying on logical structure, prototype-based approaches use exemplars as compact, interpretable representations of clusters.
  • Convex polyhedral. Convex polyhedral models represent each cluster as a convex region enclosed by a set of bounding planes. In this example, three convex regions (rectangular in shape) are formed in the $(F_{1},F_{2})$ space, each defined by the intersection of two half-planes. Samples falling within $0.5<F_{1}<2.5$ and $0.75<F_{2}<2.75$ form Cluster 1; those within $2.75<F_{1}<6$ and $2.5<F_{2}<6$ form Cluster 2; and those within $1<F_{1}<3.25$ and $1.8<F_{2}<5.5$ form Cluster 3. Interpretability arises from the explicit half-space boundaries that locally define each cluster’s convex region, while all convex polyhedra collectively establish a coherent geometric representation of the global cluster structure.
Key differences. In terms of explanation, decision-tree models yield hierarchical paths composed of sequential decisions; rule-based models provide flat logical conditions (conjunctions or disjunctions) without hierarchy; prototype-based models explain clustering through similarity to representative exemplars; and convex polyhedral models provide explanations via explicit geometric constraints formed by half-space intersections. All four approaches produce human-readable descriptions but differ in both their representational form (path, flat rule, exemplar, or hyperplane) and the aspect they make explicit, including decision order, logical structure, representativeness, and geometric boundary.
@@
(3) Interpretability level. Existing methods can be categorized into model-level and feature-level interpretability based on their degree of explainability. While most of the methods discussed in this paper focus on designing interpretable models to obtain clustering results or fitting the results of third-party algorithms, some methods also emphasize the extraction of interpretable features from complex data, or the investigation of the relationships between specific clusters and their associated features, thereby enhancing interpretability.
(4) Data modality. Finally, methods can also be classified by the nature of data they are designed to process. Common data modalities include tabular data (numerical, categorical, or mixed), sequential data (e.g., discrete sequences or time series), as well as formats such as images, text, and graphs. The appropriate interpretable model and explanatory form often depend on the structural characteristics of the underlying data.
Overall, these four criteria provide a structured taxonomy that systematizes existing interpretable clustering methods, informs future methodological development, and establishes a conceptual bridge to the taxonomies of supervised eXplainable AI (XAI).
4 Conceptual correspondence between interpretable clustering and supervised XAI
TABLE I: Conceptual correspondence between supervised explainable AI (XAI) and interpretable clustering.
Taxonomy in
supervised XAI
Meaning in
supervised learning
Correspondence in
interpretable clustering
Learning objective [14] Predict labeled outcomes with explainable reasoning faithful to both labels and features. Discover latent structures or groupings in unlabeled data while providing human-understandable rationales for the resulting clusters.
Intrinsic (ante-hoc) vs.
post-hoc
 [23]
Intrinsic models are interpretable by design (e.g., decision trees, linear models), while post-hoc methods generate explanations for already-trained black-box models. In-clustering methods embed interpretability directly into the clustering process, whereas post-clustering methods derive explanations for pre-computed cluster partitions.
Model-specific vs.
model-agnostic
 [24]
Model-specific explanations depend on the internal mechanisms or parameters of a particular model, while model-agnostic approaches can be applied to any model by analyzing its input-output behavior. All in-clustering methods are model-specific, as their interpretability stems from the internal structure of the clustering model itself, whereas post-clustering approaches can employ model-agnostic surrogate explainers to interpret black-box clustering outcomes.
Global vs.
local explanation
 [13]
Global explanations describe the overall decision logic of the model; local explanations clarify individual predictions. Cluster-level explanations correspond to global understanding of group-level patterns, whereas instance-level explanations clarify why a specific sample belongs to a cluster.
Explanation form [25] Typical forms include decision trees, rules, prototypes, feature attributions, and counterfactuals. Analogous interpretable forms include decision trees, rules, prototypes, geometric boundaries, and data-dependent descriptions.
Evaluation criteria [26] Faithfulness, fidelity, stability, simplicity, and completeness, focusing on how well explanations align with model behavior and human reasoning. Coverage and uniqueness of cluster explanations, rule or path simplicity, prototype sparsity, geometrically intuitive cluster boundaries (e.g., axis-aligned or visually interpretable regions).
Recent surveys in supervised XAI have proposed comprehensive taxonomies that categorize explanation techniques according to their position in the learning pipeline, their dependency on model internals, and the form or scope of the explanations. Representative works include Adadi and Berrada [24], Guidotti et al. [14], Gunning et al. [20], Burkart and Huber [23], and Bodria et al. [13], along with more recent efforts introducing refined evaluation frameworks and new interpretability paradigms [27, 28, 29, 30]. These studies have largely focused on predictive models, where the objective is to explain the mapping from input features to labeled outcomes. However, in unsupervised settings, no ground-truth labels exist, and the interpretive goal shifts from explaining predictions to elucidating the formation and meaning of discovered clusters. To highlight this conceptual distinction, we establish a correspondence between supervised XAI and interpretable clustering, illustrating how interpretability principles can be reformulated when supervision is absent.
The correspondence summarized in Table I demonstrates that the foundational taxonomies of supervised XAI, originally developed for labeled prediction tasks, can be extended to unsupervised learning through careful reinterpretation of objectives and evaluation principles. For example, the distinction between intrinsic and post-hoc explanations parallels the in- and post-clustering categorization discussed in Section 3. Similarly, explanation forms such as trees, rules, and prototypes have direct counterparts in interpretable clustering, as illustrated in Fig. 2.
Despite these conceptual parallels, interpretable clustering faces unique challenges that are not directly addressed in supervised XAI. The absence of ground-truth labels makes it difficult to validate explanations against reference outputs, shifting the emphasis toward internal consistency, statistical significance, and domain plausibility. These characteristics underscore the need for specialized evaluation metrics and flexible explanatory frameworks, in which in-clustering and post-clustering methods are often intertwined. Overall, this conceptual bridge illustrates how established XAI principles can inform interpretable clustering while highlighting the new methodological opportunities emerging in unsupervised interpretability research.